OpenStudy (anonymous):

In right triangle PQR where R is the right angle, PQ=13, QR=12, what is the length of PR ?

OpenStudy (anonymous):

sqrt(13^2-12^2)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

if u draw the diagram, you will find that PQ is the hypotenuse of the right angled triangle PQR. By Pythagoras Theorem, H² = P² + B² means PQ² = PR² + QR² 13² = PR² + 12² 169 = PR² + 144 169-144 = PR² 25 = PR² PR = root(25) PR = 5